# Know What You Don’t Know: Uncertainty Calibration of Process Reward Models

### v1

@@ Appendix A Extended Related Work
Recent work by Puri et al., [34] suggests that LLM reasoning processes can be framed as particle filtering—a sampling-based, probabilistic state estimation approach. Classical state estimation literature has explored methods for dynamically adjusting sample sizes based on state uncertainty [11, 12, 42, 9]. The core principle behind these adaptive sampling methods is to decrease sample sizes when uncertainty is low; this reduces computational costs while maintaining accuracy, particularly important for resource-constrained environments such as mobile robotics. Similarly, information-driven adaptive strategies have been studied in contexts of planning algorithms [17, 4, 27]. However, to the best of our knowledge, this paper represents the first exploration of instance-adaptive sampling strategies within the context of LLM inference-time scaling. We would like to remark that Yu et al., [54] also tackles the adaptive reasoning problem, and their approach shares a similar spirit with ours. Their adaptation is based on the initial success probability (i.e., question difficulty), whereas our method additionally incorporates intermediate success probabilities via the PRM. Furthermore, while they adapt the reasoning length, our method focuses on controlling the reasoning width—which we believe is a promising avenue to explore for integrating both approaches.
@@
Although state-of-the-art inference-time scaling methods frequently rely on reward models, the reliability of these models has been relatively less explored. Johnson et al., [20] introduced the theoretical framework Experts Don’t Cheat, proposing that reliable models generate predictions independent of paired hints when confident about their input; Yadkori et al., [50] further verified this idea to LLMs. Parallel research focuses on the concept of semantic entropy (SE) [23, 10], identifying hallucinations through contextual inconsistencies. Ye et al., [53] subsequently expanded SE to assess the reliability of process reward models. While these studies have made strides in quantifying LLM reliability, they do not specifically integrate reliability assessment into inference-time scaling strategies. Furthermore, rather than relying on PRM to estimate success probability (i.e., uncertainty), future work could explore uncertainty quantification tools (e.g., [13, 25, 37, 38, 33, 32]). There is also a line of calibration-oriented work on LLMs, such as Shen et al., [39], which presents another promising avenue to explore within our framework.
@@ B.1 Proof of Proposition 1
Pr[at least one $\Pr[\text{at least one success}]=1-(1-p)^{N}.$
@@ B.2 Theoretical Justifications of IAS with Beam Search
[…]
@@ B.2.1 Proof of Theorem 1
Algorithm 1 One Sided Split Conformal Quantile Regression (specialized from [35])
[…]
@@
Purely for notational convenience, let’s set $\hat{q}^{(\beta)}=\hat{r}^{(\beta)}+\delta$ where $\delta$ and $\alpha:=\tilde{\beta}$ are constants chosen using $\mathcal{I}_{1}$ such that $\delta\geq Q_{1-\tilde{\beta}}(E,\mathcal{I}_{2})$ with high probability. This step is not strictly necessary, we simply use it to avoid having to adjust our existing quantile regressor. This guarantees that $\hat{q}^{(\beta)}(X_{n+1})-Q_{1-\tilde{\beta}}(E,\mathcal{I}_{2})\geq\hat{r}^{(\beta)}(X_{n+1})$, hence
@@
$P(\text{success best-of-}N_{\mathrm{IAS}}(\hat{r}^{(\beta)}(X_{n+1},C)|X_{n+1})\geq P(\text{success best-of-}N_{\mathrm{IAS}}(P_{n+1},C)|X_{n+1})\geq C,$
@@
$P(\text{success best-of-}N_{\mathrm{IAS}}(\hat{r}^{(\beta)}(X_{n+1},C)|X_{n+1})\geq C(1-\tilde{\beta}).$
@@ C.1 Calibration Set Collection
To evaluate the calibration quality of the given LLM and PRM, we construct a validation set $\mathcal{D}_{\mathrm{val}}=\cup_{q\in\mathcal{Q}}\{(\hat{r}^{(q,i,t)},\tilde{p}^{(q,i,t)})\}$ as follows. For simplicity, we omit the superscript $q$ in the following discussion.
First, for each question $q\in\mathcal{Q}$, we generate $N_{\mathrm{val}}$ independent reasoning trajectories using the LLM: $\{\mathbf{x}^{(1)},\mathbf{x}^{(2)},\dots,\mathbf{x}^{(N_{\mathrm{val}})}\}$ where $\mathcal{Q}$ is a validation question set (i.e., green arrows in Figure 5(a)).
@@
Next, for each reasoning trajectory $\mathbf{x}^{(i)}$, and each of its prefix trajectories $x_{1:t}^{(i)}$, we generate $N_{\mathrm{MC}}$ additional follow-up trajectories conditioned on both the question and the prefix (i.e., green arrows in Figure 5(b)). By evaluating and counting the number of correct follow-up trajectories, $Z^{(i,t)}$, we estimate the reward and the success probability as: $\hat{r}^{(i,t)}=\mathsf{PRM}_{\phi}(q,x_{1:t}^{(i)})$ and $\tilde{p}^{(i,t)}\approx Z^{(i,t)}/N_{\mathrm{MC}}$ (i.e., purple arrows in Figure 5(a) and blue arrows in Figure 5(b)), respectively.
@@
Simulations are conducted with $N_{\mathrm{val}}=8$ and $N_{\mathrm{MC}}=8$, using NVIDIA V100 32GB SMX3 devices and the vLLM inference acceleration framework [24]. We adopt the standard temperature setting of $0.7$ for each LLM.
C.2 Baselines
C.2.1 Temperature Scaling
@@
C.2.2 Isotonic Regression
Isotonic regression calibration fits a non-parametric, monotonic mapping from raw scores (e.g., PRM reward) to target values (e.g., success probabilities). The method applies the Pool Adjacent Violators algorithm (PAVA) [7] to learn a nondecreasing piecewise-constant function. This mapping then stretches or compresses score regions so that, within each fitted segment, the average predicted score matches the empirical success rate. Unlike parametric methods (e.g., temperature scaling), isotonic regression makes no assumptions about the shape of the calibration curve, allowing it to flexibly adapt to arbitrary monotonic distortions in the model’s outputs.
C.2.3 Histogram Binning
Histogram binning calibration partitions the prediction interval into a fixed number of equal-width bins, then replaces each raw score with the empirical success rate of its bin. At test time, each new prediction is assigned to its corresponding interval and is binned to that interval’s mean. This simple non-parametric approach corrects systematic over- and under-confidence without assuming any particular shape of the calibration curve.
C.3 Calibration via Fine-tuning
@@
More specifically, assuming a two-token prediction scenario with logits $z_{0}$ and $z_{1}$, then softmax probabilities for two classes 00 (i.e., “bad”) and $1$ (i.e., “good”) are given by:
@@ Appendix E Additional Results
E.1 Histogram of Estimation Error
@@
As shown in Figures 6 and 7, the error densities of uncalibrated PRMs are skewed to the right, with minimal mass on the left, indicating consistent overestimation. This effect is further amplified when weaker LLMs are used to generate responses, or when evaluated on more challenging datasets such as AIME24-25.
@@
E.2 Comparison with Calibration Baselines
In this section, we present calibration results for our method and several baselines across three PRMs. In addition to reporting the average Brier score, we also plot the 95% confidence intervals. The results are shown in Figures 8, 9, and 10.
@@ E.2 Comparison with Calibration Baselines
E.3 Inference-time Scaling with IAS

### v2

@@ Appendix A Extended Related Work
Recent work by Puri et al., [37] suggests that LLM reasoning processes can be framed as particle filtering—a sampling-based, probabilistic state estimation approach. Classical state estimation literature has explored methods for dynamically adjusting sample sizes based on state uncertainty [12, 13, 45, 10]. The core principle behind these adaptive sampling methods is to decrease sample sizes when uncertainty is low; this reduces computational costs while maintaining accuracy, particularly important for resource-constrained environments such as mobile robotics. Similarly, information-driven adaptive strategies have been studied in contexts of planning algorithms [18, 5, 29]. However, to the best of our knowledge, this paper represents the first exploration of instance-adaptive sampling strategies within the context of LLM inference-time scaling.
We would like to remark that Yu et al., [59] also tackles the adaptive reasoning problem, and their approach shares a similar spirit with ours. Their adaptation, however, is based on the initial success probability (i.e., question difficulty), whereas our method additionally adapts at intermediate reasoning steps (and thus is extendable to beam search). Specifically, our approach is enabled by calibrating PRMs, which allows us to estimate success probabilities throughout the multi-step reasoning process—a direction that has not been explored before. Furthermore, while they adapt the reasoning length, our method focuses on controlling the reasoning width (i.e., sample size), which we believe is a promising avenue to explore for integrating both approaches.
@@
Although state-of-the-art inference-time scaling methods frequently rely on reward models, the reliability of these models has been relatively less explored. Johnson et al., [21] introduced the theoretical framework Experts Don’t Cheat, proposing that reliable models generate predictions independent of paired hints when confident about their input; Yadkori et al., [54] further verified this idea extends to LLMs. Parallel research focuses on the concept of semantic entropy (SE) [25, 11], identifying hallucinations through contextual inconsistencies. Ye et al., [58] subsequently expanded SE to assess the reliability of process reward models. While these studies have made strides in quantifying LLM reliability, they do not specifically integrate reliability assessment into inference-time scaling strategies. Furthermore, rather than relying on PRM to estimate success probability (i.e., uncertainty), future work could explore uncertainty quantification tools (e.g., [14, 27, 40, 41, 36, 35]). There is also a line of calibration-oriented work on LLMs, such as Shen et al., [42], which presents another promising avenue to explore within our framework.
@@ B.1 Proof of Proposition 1
$\Pr[\text{at least one success}]=1-(1-p)^{N}.$
@@ B.2 Theoretical Justifications of IAS with Beam Search
$K_{\mathrm{IAS}}\;=\;\min\bigl\{k\mid\;N_{\mathrm{IAS}}\!\bigl(p_{k},C\bigr)\;\leq\;k\times M\;,\;1\leq k\leq K_{\max}\bigr\}.$
@@ B.2.1 Proof of Theorem 1
Algorithm 1 One-Sided Split Conformal Quantile Regression (specialized from [38])
1:Data $(X_{i},Y_{i})\in\mathbb{R}^{p}\times\mathbb{R},\ 1\leq i\leq n$; Miscoverage level $\alpha\in(0,1)$; Quantile regression algorithm $\mathcal{A}$ (e.g. the QR method described in the main text [^3] The theorem works for any quantile estimation method.).
@@
Purely for notational convenience, let’s set $\hat{q}^{(\beta)}=\hat{r}^{(\beta)}+\delta$ where $\delta$ and $\alpha:=\tilde{\beta}$ are constants chosen using $\mathcal{I}_{1}$ such that $\delta\geq Q_{1-\tilde{\beta}}(E,\mathcal{I}_{2})$ with high probability. This step is not strictly necessary; we simply use it to avoid having to adjust our existing quantile regressor. This guarantees that $\hat{q}^{(\beta)}(X_{n+1})-Q_{1-\tilde{\beta}}(E,\mathcal{I}_{2})\geq\hat{r}^{(\beta)}(X_{n+1})$, hence
@@
[…]
@@
[…]
@@ C.1 Calibration Set Collection
To evaluate the calibration quality of a given LLM and PRM pair, we construct a validation set $\mathcal{D}_{\mathrm{val}}=\cup_{q\in\mathcal{Q}_{\mathrm{val}}}\{(\hat{r}^{(q,i,t)},\tilde{p}^{(q,i,t)})\}$ through a systematic three-stage process, as described in Section 3.1 of the main text. Here we provide additional implementation details and statistical analysis.
Data Collection Procedure.
For notational simplicity, we omit the superscript $q$ in the following discussion, with the understanding that all operations are performed independently for each question $q\in\mathcal{Q}_{\mathrm{val}}$.
@@
For each question $q\in\mathcal{Q}_{\mathrm{val}}$, we first generate $N_{\mathrm{val}}=8$ independent reasoning trajectories using the target LLM: $\{\mathbf{x}^{(1)},\mathbf{x}^{(2)},\dots,\mathbf{x}^{(N_{\mathrm{val}})}\}$ (illustrated as green arrows in Figure 5(a)). Next, for each reasoning trajectory $\mathbf{x}^{(i)}$ and each of its prefix trajectories $x_{0:t}^{(i)}$ where $t\in\{0,1,\dots,T^{(i)}\}$, we generate $N_{\mathrm{MC}}=8$ additional follow-up trajectories conditioned on both the question and the prefix (green arrows in Figure 5(b)). By evaluating and counting the number of correct follow-up trajectories $Z^{(i,t)}$, we estimate the PRM’s predicted reward $\hat{r}^{(i,t)}=\mathsf{PRM}_{\phi}(q,x_{0:t}^{(i)})$ and the empirical success probability $\tilde{p}^{(i,t)}\approx Z^{(i,t)}/N_{\mathrm{MC}}$ (purple arrows in Figure 5(a) and blue arrows in Figure 5(b)).
Hardware and Software Configuration.
@@
All simulations are conducted using NVIDIA V100 32GB SMX3 devices with the vLLM inference acceleration framework [26]. We employ the standard sampling configuration: top_p = 1.0, top_k = -1 (considering all tokens in the vocabulary), and temperature $T=0.7$ for all LLMs, consistent with best practices in recent reasoning model studies [2, 37].
C.2 Statistical Properties and Label Quality
Unbiased Estimation.
The per-prefix Monte Carlo estimator $\tilde{p}^{(i,t)}$ is statistically unbiased for the true success probability. While individual estimates with $N_{\mathrm{MC}}=8$ have non-negligible variance ($\sigma_{\varepsilon}^{2}=p(1-p)/N_{\mathrm{MC}}\approx 0.02$ for typical $p\approx 0.8$), our approach compensates through large-scale aggregation.
Aggregate Accuracy.
We collect between M=40, [^0] M=40{,}000 and 80,00080{,}000 prefix-label pairs per PRM–LLM combination. Aggregating over this large number of samples substantially reduces the standard error of our calibration metrics. Specifically, the standard error of the Brier score (our primary calibration measure) scales approximately as:
$\text{SE}\approx\sqrt{\frac{2\sigma_{\varepsilon}^{4}+4\sigma_{\varepsilon}^{2}\cdot\mathrm{MSE}_{\text{true}}}{M}},$ (3)
where $\sigma_{\varepsilon}^{2}=\frac{p(1-p)}{N_{\mathrm{MC}}}\approx 0.02$ represents the per-instance variance (assuming typical success probability $p\approx 0.8$), and $\mathrm{MSE}_{\text{true}}\approx 0.08$ represents the model’s intrinsic prediction error. With our collection size of $M\geq 40{,}000$, this standard error becomes sufficiently small that our reported calibration metrics are statistically reliable, and the influence of per-instance noise on global trends is negligible.
Impact on Quantile Regression.
While the aggregate statistics are accurate, the instance-wise standard error does artificially inflate the variance of the target distribution in our quantile regression objective. Increasing $N_{\mathrm{MC}}$ could potentially improve our method’s performance by providing cleaner training signals. However, our experimental results demonstrate that our calibration approach is effective even with $N_{\mathrm{MC}}=8$, achieving substantial improvements in calibration metrics across all evaluated PRM–LLM pairs.
C.3 Additional Implementation Details
Correctness Evaluation.
[…]
PRM Reward Calculation.
The reward computation follows the official implementation provided by each PRM developer. In our experiments, we extract the final score predicted at the last reasoning step, as this approach has been shown to perform well in inference-time scaling regimes [2, 43].
For Qwen-PRM, we use the following prompt format. The system prompt is:
<|im_start|>system
[…]
The user prompt format is:
<|im_start|>user
\${Question}<|im_end|>
<|im_start|>assistant
\${Prefix}<extra_0><|im_end|><|endoftext|>
where \$Question represents the problem statement and \$Prefix denotes the reasoning trajectory prefix (which is an empty string in the case when applying IAS for the Best-of-N regime). The reward score is computed based on the token probability at the special <extra_0> marker, which corresponds to the final reasoning step. Similar prompt structures are used for other PRMs, following their respective documentation.
Data and Code Release.
To ensure full reproducibility and facilitate future research, we publicly release our complete codebase, including all prompt templates, sampling configurations, and evaluation scripts. Our calibration datasets—comprising $\bigl(q,x_{0:t}^{(q,i)},\tilde{p}^{(q,i,t)}\bigr)$ triplets of question, prefix trajectory, and empirical success probability—are available for multiple reasoning LLMs at https://huggingface.co/datasets/young-j-park/prm_calibration. The dataset includes calibration data for all 18 PRM–LLM pairs evaluated in this work.
C.4 Baselines
C.4.1 Temperature Scaling
@@
C.4.2 Isotonic Regression
Isotonic regression calibration fits a non‐parametric, monotonic mapping from raw scores (e.g., PRM reward) to target values (e.g., success probabilities). The method applies the Pool Adjacent Violators algorithm (PAVA) [8] to learn a nondecreasing piecewise‐constant function. This mapping then stretches or compresses score regions so that, within each fitted segment, the average predicted score matches the empirical success rate. Unlike parametric methods (e.g., temperature scaling), isotonic regression makes no assumptions about the shape of the calibration curve, allowing it to flexibly adapt to arbitrary monotonic distortions in the model’s outputs.
C.4.3 Histogram Binning
Histogram binning calibration partitions the prediction interval into a fixed number of equal-width bins, then replaces each raw score with the empirical success rate of its bin. At test time, each new prediction is assigned to its corresponding interval and is binned to that interval’s mean. This simple non‐parametric approach corrects systematic over‐ and under‐confidence without assuming any particular shape of the calibration curve.
C.5 Calibration via Fine-tuning
@@
More specifically, assuming a two-token prediction scenario with logits $z_{0}$ and $z_{1}$, then softmax probabilities for two classes 0 (i.e., “bad”) and $1$ (i.e., “good”) are given by:
@@
C.6 Model Calibration
Calibration error metrics assess whether a model’s predicted probabilities are consistent with observed accuracy. Two widely used calibration metrics are the Brier score [3] and the expected calibration error (ECE) [32]. Brier score quantifies the mean-squared difference between predicted probabilities $\hat{y}^{(l)}$ and ground truth labels $y^{(l)}$ over $N$ instances, where $l=1,\dots,N$:
$\text{Brier}\;=\;\frac{1}{N}\,\sum_{l=1}^{N}\bigl(\hat{y}^{(l)}-y^{(l)}\bigr)^{2}.$
$\text{ECE}\;=\;\sum_{b=1}^{B}\frac{|\mathcal{I}_{b}|}{N}\,\bigl|\operatorname{conf}(\mathcal{B}_{b})-\operatorname{acc}(\mathcal{B}_{b})\bigr|,$
where $\operatorname{conf}(\mathcal{B}_{b})=\tfrac{1}{|\mathcal{I}_{b}|}\sum_{l\in\mathcal{I}_{b}}\hat{y}^{(l)}$ is the average confidence in bin $b$ and $\operatorname{acc}(\mathcal{B}_{b})=\tfrac{1}{|\mathcal{I}_{b}|}\sum_{l\in\mathcal{I}_{b}}y^{(l)}$ is the empirical accuracy. Variants of ECE include adaptive calibration error (AdaptiveCE) [34] and average calibration error (AverageCE) [33], which uses adaptive bin intervals and equal weighting across bins, respectivley.
@@ Appendix E Additional Results
E.1 Calibration Analysis of Larger-Scale PRMs
We extend our calibration analysis to larger-scale process reward models by evaluating the 72B parameter variant on both MATH500 and AIME24-25 datasets. This analysis addresses whether the calibration challenges observed in 7B models persist at larger scales, and whether our calibration method remains effective across different model sizes.
Table 4 and Figure 6 present a comprehensive comparison of calibration performance across three configurations: uncalibrated 7B models, uncalibrated 72B models, and our calibrated 7B models. We evaluate six different base language models across two datasets, reporting both Brier score and Positive Brier score metrics for each configuration.
The results demonstrate that while larger 72B models exhibit reduced overestimation compared to their 7B counterparts, they still suffer from significant miscalibration issues. This observation is expected, as both model were trained with same policy model.
These results validate two key insights: (1) the calibration problem persists across model scales and cannot be solved simply by increasing model capacity, and (2) explicit calibration methods can effectively address miscalibration. Notably, our calibrated 7B model not only outperforms the uncalibrated 7B baseline but also surpasses the much larger uncalibrated 72B model, demonstrating that targeted calibration is more effective than naive scaling.
Table 4: Calibration performance comparison across model scales on MATH500 and AIME24-25 datasets (lower is better). Best results for each metric within each row are shown in bold. Results demonstrate that explicit calibration of Qwen-PRM-7B models outperforms even uncalibrated Qwen-PRM-72B models, highlighting the importance of targeted calibration over naive model scaling.
Dataset Model Uncalibrated 7B Uncalibrated 72B Calibrated 7B
Brier Pos. Brier Brier Pos. Brier Brier Pos. Brier
MATH500 Llama-3.2-1B 0.241 0.223 0.211 0.078 0.069 0.047
Llama-3.1-8B 0.205 0.177 0.177 0.111 0.121 0.099
Qwen-2.5-1.5B 0.154 0.130 0.155 0.106 0.127 0.107
Qwen-2.5-7B 0.101 0.085 0.118 0.057 0.082 0.053
R1-Llama-8B 0.161 0.114 0.193 0.072 0.089 0.055
R1-Qwen-7B 0.148 0.106 0.176 0.075 0.083 0.058
AIME24-25 Llama-3.2-1B 0.194 0.192 0.157 0.063 0.003 0.001
Llama-3.1-8B 0.227 0.223 0.183 0.122 0.041 0.035
Qwen-2.5-1.5B 0.330 0.322 0.256 0.176 0.073 0.053
Qwen-2.5-7B 0.289 0.282 0.205 0.139 0.072 0.066
R1-Llama-8B 0.385 0.371 0.237 0.204 0.078 0.030
R1-Qwen-7B 0.414 0.402 0.260 0.215 0.069 0.034
(a) Uncalibrated Qwen-PRM-72B
(b) Uncalibrated Qwen-PRM-72B
Figure 6: Histogram of signed deviation (i.e., estimation error) for Qwen-PRM-72B on the MATH500 (in-distribution) and AIME24-25 (out-of-distribution) datasets. Positive error indicates overestimation. While larger 72B model exhibit reduced overestimation compared to their 7B counterparts, they still suffer from significant miscalibration issues.
E.2 Histogram of Estimation Error
@@
As shown in Figures 7 and 8, the error densities of uncalibrated PRMs are skewed to the right, with minimal mass on the left, indicating consistent overestimation. This effect is further amplified when weaker LLMs are used to generate responses, or when evaluated on more challenging datasets such as AIME24-25.
@@
E.3 Comparison with Calibration Baselines
In this section, we present calibration results for our method and several baselines across three PRMs. In addition to reporting the average Brier score, we also plot the 95% confidence intervals. The results are shown in Figures 9, 10, and 11.
@@ E.3 Comparison with Calibration Baselines
We additionally compare our method against Adaptive Temperature Scaling (ATS) [53, 42], a recent calibration method that learns instance-specific temperature parameters. Since ATS was originally designed for standard language model outputs rather than PRM regression-style scores, we adapted it for our setting. Specifically, we modified ATS to train a transformer head that dynamically selects temperature parameters based on the PRM-generated hidden states of both the question and solution prefix, enabling context-aware calibration similar to our approach.
Table 5 presents comprehensive comparisons across three calibration metrics. The modified ATS consistently outperforms simpler baselines, validating the importance of context-aware calibration. On the in-distribution MATH500 dataset, ATS achieves competitive performance—for instance, reducing AdaCE from 0.241 to 0.178 for R1-Qwen-7B. However, our method achieves superior calibration with AdaCE of 0.107 for the same model.
The performance gap becomes more pronounced on the out-of-distribution AIME24-25 dataset. While ATS provides improvements over uncalibrated models (e.g., reducing AdaCE from 0.558 to 0.440 for R1-Qwen-7B), our method demonstrates substantially better generalization with an AdaCE of only 0.090. This pattern holds consistently across all models and metrics. For example, on Llama-3.2-1B, our method achieves an AdaCE of 0.011 compared to 0.125 for ATS, and a Brier score of 0.003 compared to 0.065.
Table 5: Calibration error for various methods and models on MATH500 (in-distribution) and AIME24-25 (out-of-distribution). Values are presented as adaptive calibration error (AdaCE) and Brier scores, where lower is better. The best result for each metric within a row is bolded. We use Qwen-PRM-7B. Acronyms: TS (Temperature Scaling), ATS (Adpative Temperature Scaling), IR (Isotonic Regression), and HB (Histogram Binning).
Dataset Model Uncalibrated TS ATS IR HB Calibrated (Ours)
AdaCE Brier AdaCE Brier AdaCE Brier AdaCE Brier AdaCE Brier AdaCE Brier
MATH500 Llama-3.2-1B .283 .241 .265 .190 .181 .099 .212 .097 .213 .097 .081 .069
Llama-3.1-8B .263 .205 .244 .168 .237 .130 .321 .166 .323 .166 .167 .121
Qwen-2.5-1.5B .218 .154 .205 .140 .205 .123 .271 .145 .274 .146 .155 .127
Qwen-2.5-7B .146 .101 .135 .093 .135 .076 .201 .092 .205 .094 .100 .082
R1-Llama-8B .251 .161 .223 .149 .196 .107 .297 .141 .298 .141 .113 .089
R1-Qwen-7B .241 .148 .202 .140 .178 .103 .289 .132 .289 .132 .107 .083
AIME24-25 Llama-3.2-1B .236 .194 .210 .193 .125 .065 .177 .065 .177 .065 .011 .003
Llama-3.1-8B .284 .227 .248 .206 .281 .125 .392 .215 .396 .215 .086 .041
Qwen-2.5-1.5B .401 .330 .393 .293 .367 .184 .477 .331 .486 .337 .105 .073
Qwen-2.5-7B .355 .289 .354 .249 .271 .146 .390 .250 .404 .261 .098 .072
R1-Llama-8B .526 .385 .528 .310 .438 .220 .590 .400 .588 .399 .128 .078
R1-Qwen-7B .558 .414 .570 .340 .440 .230 .613 .430 .613 .430 .090 .069
E.4 Inference-time Scaling with IAS